# Compute Optimal Inference and Provable Amortisation Gap in Sparse Autoencoders

### v1

@@ Appendix C Decoder weight analysis
A useful method for gaining insight into the behavior of our models is through examining the final weights of the decoder. Specifically, we visualize $W^{\top}W$, an $N\times N$ matrix, for three scenarios: when $N$ equals the true sparse dimensionality, when $N$ exceeds it, and when $N$ is smaller than the true dimensionality.
@@ Appendix D MLP Ablations
Appendix E Including a bias parameter
@@
However, Theorem 1 doesn’t rely on having biases, and although it generalises to the case with biases, we would like to be able to simplify our study by not including them. Thus, we show in Figure 13 that biases have no statistically significant effect on reconstruction loss, latent MCC, dictionary MCC, or L0, for any of the models, except for the L0 and MCC of the MLP, which achieves a higher MCC without bias at the cost of a greater L0.
@@
Appendix F Comparison with traditional dictionary learning methods
@@
F.1 Optimised Sparse Autoencoders and Sparse Coding
@@
F.1.1 Advanced Sparse Autoencoder Techniques
@@
F.1.2 Optimised Sparse Coding Approaches
@@ F.2 Top-k sparse coding
Appendix G Measuring FLOPs
@@ G.1 Sparse Coding
$\text{FLOPs}_{\text{SC-train}}=n_{\text{eff}}\cdot(\text{FLOPs}_{\text{forward}}+\text{FLOPs}_{\text{loss}}+\text{FLOPs}_{\text{backward}}+\text{FLOPs}_{\text{update}})$ (18)
@@
$\text{FLOPs}_{\text{forward}}$ $=\text{FLOPs}_{\text{SC-inf}}$ (19)
$\text{FLOPs}_{\text{loss}}$ $=2Mn_{b}+Nn_{b}$ (20)
$\text{FLOPs}_{\text{backward}}$ $\approx 2\cdot\text{FLOPs}_{\text{forward}}$ (21)
$\text{FLOPs}_{\text{update}}$ =$=\begin{cases}Nn_{b}+MN&\text{if learning }D\\ Nn_{b}&\text{otherwise}\end{cases}$ (22)
G.2 Sparse Autoencoder (SAE)
@@
$\text{FLOPs}_{\text{SAE-train}}=n_{\text{eff}}\cdot(\text{FLOPs}_{\text{forward}}+\text{FLOPs}_{\text{backward}})$ (23)
@@
$\text{FLOPs}_{\text{forward}}$ =$=\begin{cases}5MN+N&\text{if learning }D\\ 4MN+N&\text{otherwise}\end{cases}$ (24)
$\text{FLOPs}_{\text{backward}}$ $=N+(2NM+N)+2NM+2(MN+N)+\begin{cases}2NM&\text{if learning }D\\ 0&\text{otherwise}\end{cases}$ (25)
@@
$\text{FLOPs}_{\text{SAE-inf}}=(4MN+N)\cdot n_{s}$ (26)
G.3 Multilayer Perceptron (MLP)
@@
$\text{FLOPs}_{\text{MLP-train}}=n_{\text{eff}}\cdot(\text{FLOPs}_{\text{forward}}+\text{FLOPs}_{\text{backward}})$ (27)
@@
$\text{FLOPs}_{\text{forward}}$ =$=\begin{cases}2MH+H+2HN+N+2NM+MN&\text{if learning }D\\ 2MH+H+2HN+N+2NM&\text{otherwise}\end{cases}$ (28)
$\text{FLOPs}_{\text{backward}}$ $=N+(2NH+N)+H+(2MH+H)+2NM+2(MH+H+HN+N)$ (29)
@@
$\text{FLOPs}_{\text{MLP-inf}}=(2MH+H+2HN+N+2NM)\cdot n_{s}$ (30)
G.4 SAE with Inference-Time Optimisation (SAE+ITO)
For SAE+ITO, we calculate the additional FLOPs required for optimizing the codes during inference:
$\text{FLOPs}_{\text{ITO}}=(MN+N+n_{\text{iter}}\cdot(4MN+2M+11N))\cdot n_{s}$ (31)
@@ G.4 SAE with Inference-Time Optimisation (SAE+ITO)
These FLOP calculations provide a consistent measure of computational cost across all methods, allowing for fair comparisons of efficiency and performance trade-offs.
Appendix H Automated interpretability
In this section, we describe the automated interpretability pipeline used to understand and evaluate the features learned by sparse autoencoders (SAEs) and other models in the context of neuron activations within large language models (LLMs). The pipeline consists of two distinct tasks: feature interpretation and feature scoring. These tasks allow us to generate hypotheses about individual feature activations and to determine whether specific features are likely to activate given particular token contexts.
H.1 Feature Interpreter Prompt
@@
H.2 Feature Scorer Prompt
@@
H.3 Evaluation of Automated Interpretability

### v2

@@ Appendix E Decoder weight analysis
A useful method for gaining insight into the behaviour of our models is through examining the final weights of the decoder. Specifically, we visualise $W^{\top}W$, an $N\times N$ matrix, for three scenarios: when $N$ equals the true sparse dimensionality, when $N$ exceeds it, and when $N$ is smaller than the true dimensionality.
@@ Appendix F MLP Ablations
Appendix G Including a bias parameter
@@
However, Theorem 3.1 doesn’t rely on having biases, and although it generalises to the case with biases, we would like to be able to simplify our study by not including them. Thus, we show in Figure 16 that biases have no statistically significant effect on reconstruction loss, latent MCC, dictionary MCC, or L0, for any of the models, except for the L0 and MCC of the MLP, which achieves a higher MCC without bias at the cost of a greater L0.
@@
Appendix H Comparison with traditional dictionary learning methods
@@
H.1 Optimised Sparse Autoencoders and Sparse Coding
@@
H.1.1 Advanced Sparse Autoencoder Techniques
@@
H.1.2 Optimised Sparse Coding Approaches
@@ H.2 Top-k sparse coding
Appendix I Measuring FLOPs
@@ I.1 Sparse Coding
$\text{FLOPs}_{\text{SC-train}}=n_{\text{eff}}\cdot(\text{FLOPs}_{\text{forward}}+\text{FLOPs}_{\text{loss}}+\text{FLOPs}_{\text{backward}}+\text{FLOPs}_{\text{update}})$
@@
$\text{FLOPs}_{\text{forward}}$ $=\text{FLOPs}_{\text{SC-inf}}$
$\text{FLOPs}_{\text{loss}}$ $=2Mn_{b}+Nn_{b}$
$\text{FLOPs}_{\text{backward}}$ $\approx 2\cdot\text{FLOPs}_{\text{forward}}$
$\text{FLOPs}_{\text{update}}$ =$=\begin{cases}Nn_{b}+MN&\text{if learning }D\\ Nn_{b}&\text{otherwise}\end{cases}$
I.2 Sparse Autoencoder (SAE)
@@
$\text{FLOPs}_{\text{SAE-train}}=n_{\text{eff}}\cdot(\text{FLOPs}_{\text{forward}}+\text{FLOPs}_{\text{backward}})$
@@
$\text{FLOPs}_{\text{forward}}$ =$=\begin{cases}5MN+N&\text{if learning }D\\ 4MN+N&\text{otherwise}\end{cases}$
$\text{FLOPs}_{\text{backward}}$ $=N+(2NM+N)+2NM+$
$2(MN+N)+\begin{cases}2NM&\text{if learning }D\\ 0&\text{otherwise}\end{cases}$
@@
$\text{FLOPs}_{\text{SAE-inf}}=(4MN+N)\cdot n_{s}$
I.3 Multilayer Perceptron (MLP)
@@
$\text{FLOPs}_{\text{MLP-train}}=n_{\text{eff}}\cdot(\text{FLOPs}_{\text{forward}}+\text{FLOPs}_{\text{backward}})$
@@
$\text{FLOPs}_{\text{MLP-inf}}=(2MH+H+2HN+N+2NM)\cdot n_{s}$
I.4 SAE with Inference-Time Optimisation (SAE+ITO)
For SAE+ITO, we calculate the additional FLOPs required for optimising the codes during inference:
$\text{FLOPs}_{\text{ITO}}=(MN+N+n_{\text{iter}}\cdot(4MN+2M+11N))\cdot n_{s}$
@@ I.4 SAE with Inference-Time Optimisation (SAE+ITO)
Appendix J Automated interpretability
Table 1: Example interpretations from MLP and SAE neurons, shown with their F1 scores.
Model Interpretation F1 Score
MLP Activates on the token “to” when used to introduce an infinitive verb indicating purpose or intent, promoting verbs that express actions or goals 0.899
Activates on concrete and functional nouns or specific actions that are often part of a list or enumeration 0.899
Activates on the token “than” as part of a comparative structure, aiding in predicting terms used for comparison or establishing norms 1.000
SAE Activates on parentheses and colons used in structured timestamps, date-time formats, and categorisation notations 1.000
Activates on tokens within contexts related to font and text styling options, typically presented in a technical or settings menu format 1.000
Activates on the token “first” within the formulaic expression “first come, first served basis” 1.000
In this section, we describe the automated interpretability pipeline used to understand and evaluate the features learned by sparse autoencoders (SAEs) and other models in the context of neuron activations within large language models (LLMs). The pipeline consists of two tasks: feature interpretation and feature scoring. These tasks allow us to generate hypotheses about individual feature activations and to determine whether specific features are likely to activate given particular token contexts.
J.1 Feature Interpreter Prompt
@@
J.2 Feature Scorer Prompt
@@
J.3 Evaluation of Automated Interpretability
@@
Table 1 presents illustrative examples of interpretations from both MLP and SAE neurons, showing how our automated pipeline can identify specific linguistic patterns and assign quantitative reliability scores.